Californium, Cf (atomic mass 251, atomic number 98), is the most expensive
element. It was sold for $10.00 per μg in 1992. (1μg = 10−6g = 10−9kg)
[a]How much would 1kg of Cf cost?
[b]How many electrons are in 1μg of Cf?
[c]How much is that per electron? (This may be a small number.)
[d]Are electrons the ultimate cheap gift?
Answer #1

mass of 1 mole=251 g

cost of 1 kg =10*10^9 =10^8 dollars

let electron present in 1 ug=x

no of moles in 1 ug=(1/251)*10^(-6)=3.984*10^(-9) moles

1 mol=6.023*10^23 atom

so total atom of cf present=2.399*10^15 atom

1 atom=98 electron

so total electron=2.35*10^(17) electron

(c)cost per electron=cost/no of electron=$4.25*10^(-17) per electron

(d)Yes